# Open-World Task and Motion Planning via Vision-Language Model Inferred Constraints

### v2

@@ A-G Code as Policies Full Prompt
1You are operating in an environment that has access to the following classes: {env_structs}.
2
@@
3You have access to the following variables. Note that TABLE_AABB.lower and TABLE_AABB.upper are tuples of length 3 that represent the lower and upper bounds on the x, y, and z positions that are on the table.
4{typed_variables}
5
6You also have access to the following objects in the environment.
7Each of these has type RavenObject‘. You are not able to use other objects:
8{object_names}
9
@@
10Note that when you use/invoke these objects, do not use them as a string (i.e., do not provide quotes ’’). Instead
@@
11use them directly (e.g. use banana, not ’banana’). Use the name of the object as a string only where explicitly required
12(e.g. by the GraspSampler).
13
@@
14Here’s the poses of all the objects in the initial state of the scene (depicted in an attached image). Note that the axes of each object is shown as red (x-axis), green (y-axis), and blue (z-axis).
@@
15Note that roll is rotation about the x-axis, pitch is rotation about the y-axis and yaw is rotation about the z-axis. Pay careful attention to the axes and the current orientations of objects in the initial state as and when you decide to write functions involving orientations.
16{init_state}
17
18
@@
19You also have access to helper functions whose signatures and docstrings are shown below. Pay careful attention to the arguments and return values of each function.
20Helper function signatures:
21{helper_functions_and_docstrings}
22
@@
23You have access to the following set of skills expressed as pddl predicates followed by descriptions.
@@
24You have no other skills you can use, and you must exactly follow the number of inputs described below.
@@
25The coordinate axes are x, y, z where x is distance from the robot base, y is left/right from the robot base, and z is the height off the table.
26
27Action("pick", [[o], g])
@@
28Pick up object o at grasp g sampled from a grasp sampler. Note that you should use the object o directly and not use its name (e.g. banana and not banana’).
29
30Action("place_ontop", [[o, s], p])
@@
31If holding an object o (e.g. ‘banana‘) place the object ontop surface s (e.g. ‘table‘) at pose p. Note that you should invoke s and o as objects directly and not use their names (e.g. banana and not banana’).
32
33Action("place_inside", [[o, c], p])
@@
34If holding an object o (e.g. ‘banana‘) place the object inside container c (e.g. ‘bowl‘) at placement pose p. Note that you should invoke o and c as objects directly and not use their names (e.g. banana and not banana’).
35
36Action("pour", [[o, c], p])
@@
37If holding an object ‘o‘ (e.g. ‘tomato_soup_can‘) pour from the object onto/inside of container ‘c‘ (e.g. ‘bowl‘) at a pouring pose dictated by p. Specifically, p is 4 numbers: <x, y, z> corresponding to the position to reach, and <theta> corresponding to the angle (in radians; -3.14 to 3.14) to tilt the hand to accomplish the pour. Note that you should invoke o and c as objects directly and not use their names (e.g. banana and not banana’).
38
@@
39Finally, to help you generate continuous parameters required for these skills, you have access to the following samplers.
40
41@dataclass
42class GraspSampler(Sampler):
43 curr_state: RavenState
44 env: RavenYCBEnv
45 object_name: str
46
47 def sample(self, rng: np.random.Generator) -> List[float]:
48 """Simply return some value within the AABB of the object, and at any
49 orientation."""
50 aabb_lb, aabb_ub = get_aabb_bounds(
51 self.curr_state, self.env, self.object_name
52 )
53 lb = (aabb_lb[0], aabb_lb[1], aabb_lb[2], -3.14159265, -3.14159265, -3.14159265)
54 ub = (aabb_ub[0], aabb_ub[1], aabb_ub[2], 3.14159265, 3.14159265, 3.14159265)
55 return rng.uniform(lb, ub).tolist()
56
57
58@dataclass
59class PlaceSampler(Sampler):
60 curr_state: RavenState
61 env: RavenYCBEnv
62 object_name_to_place_on_or_inside: str
63
64 def sample(self, rng: np.random.Generator) -> List[float]:
65 """Simply return some value within the AABB of the object, and at any
66 orientation."""
67 min_drop_height = 0.01
68 max_drop_height = 0.35
69 aabb_lb, aabb_ub = get_aabb_bounds(
70 self.curr_state, self.env,
71 self.object_name_to_ place_on_or_inside
72 )
73 lb = (
74 aabb_lb[0],
75 aabb_lb[1],
76 aabb_lb[2] + min_drop_height,
77 -3.14159265,
78 -3.14159265,
79 -3.14159265,
80 )
81 ub = (
82 aabb_ub[0],
83 aabb_ub[1],
84 aabb_ub[2] + max_drop_height,
85 3.14159265,
86 3.14159265,
87 3.14159265,
88 )
89 return rng.uniform(lb, ub).tolist()
90
91
92@dataclass
93class PourSampler(Sampler):
94 curr_state: RavenState
95 env: RavenYCBEnv
96 holding_obj_name: str
@@
97 obj_name_to_pour_into_or_ontop: str
98
99 def sample(self, rng: np.random.Generator) -> List[float]:
100 """Return a position sampled from above obj_name_to_pour_ into_or_ontop
101 and a pitch randomly sampled between bounds."""
102 # We sample (x, y, z, pitch.)
103 obj_holding_lb, obj_holding_ub = get_aabb_bounds(
104 self.curr_state, self.env, self.holding_obj_name
105 )
106 min_pour_height = obj_holding_ub[2] - obj_holding_lb[2]
107 max_pour_height = min_pour_height * 2
108 aabb_lb, aabb_ub = get_aabb_bounds(
109 self.curr_state, self.env, self.obj_name_to_pour _into_or_ontop
110 )
111 pitch_range = (-3.14, 3.14)
112 lb = (aabb_lb[0], aabb_lb[1], aabb_lb[2] + min_pour_height, pitch_range[0])
113 ub = (aabb_ub[0], aabb_ub[1], aabb_ub[2] + max_pour_height, pitch_range[1])
114 return rng.uniform(lb, ub).tolist()
115
@@
116Your goal is to generate a python function that returns a plan that performs the provided task. This function can
@@
117use helper functions that must be defined within the scope of the function itself.
118
@@
119The main function should be named EXACTLY ‘gen_plan‘, and it should take in only one parameter corresponding to the environment state as input. Do not change the names. Do not create any additional classes or overwrite any existing ones. You are only allowed to create helper functions inside the ‘gen_plan‘ function.
120
121Current provided goal: {provided_task_goal}
122
123Here is an example ‘gen_plan‘ function for a different task:
124#define user
125Init state:
126bowl: Pose=RavenPose(x=-0.09269248694181442, y=-0.7042829990386963, z=0.026169249787926674, roll=0.0, pitch=-0.0, yaw=0.8605557025412023)
127banana: Pose=RavenPose(x=0.17416073374449514, y=-0.33348321026557554, z=0.02017684663429707, roll=5.081222700168695e-05, pitch=0.00013538346655467005, yaw=-3.0371082921616765)
128plate: Pose=RavenPose(x=-0.11636300384998322, y=-0.4429782032966614, z=0.014744692512349077, roll=7.884650441866775e-28, pitch=-7.554679105908491e-28, yaw=2.245637386214381)
129table: Pose=RavenPose(x=0.0, y=-0.5, z=0.0, roll=0.0, pitch=-0.0, yaw=0.0)
130Task goal: put the lemon on the plate and the banana on the table
131
132#define assistant
133‘‘‘python
134def gen_plan(initial:RavenState):
135 plan = []
136 place_pose = PlaceSampler(initial, env, "plate").sample(rng)
137 lemon_grasp = GraspSampler(initial, env, "lemon").sample(rng)
138 plan.append(Action("pick", [[lemon], lemon_grasp]))
@@
139 plan.append(Action("place_ontop", [[lemon, plate], place_pose]))
140 place_pose = PlaceSampler(initial, env, "table").sample(rng)
141 banana_grasp = GraspSampler(initial, env, "banana").sample(rng)
142 plan.append(Action("pick", [[banana], banana_grasp]))
@@
143 plan.append(Action("place_ontop", [[banana, table], RavenPose(x=x_place, y=y_place, z=0.02)]))
144 return plan
145‘‘‘
146
@@
147Here is another example for a different goal involving the same objects and initial state. Note that in this case, the bowl is blue.
@@
148Also note that the banana needs to be rotated so that it fits into the bowl; in general you should pay careful
@@
149attention to any angular constraints that might be important for solving different tasks.
150
151#define user
152Init state:
153bowl: Pose=RavenPose(x=-0.09269248694181442, y=-0.7042829990386963, z=0.026169249787926674, roll=0.0, pitch=-0.0, yaw=0.8605557025412023)
154banana: Pose=RavenPose(x=0.17416073374449514, y=-0.33348321026557554, z=0.02017684663429707, roll=5.081222700168695e-05, pitch=0.00013538346655467005, yaw=-3.0371082921616765)
155plate: Pose=RavenPose(x=-0.11636300384998322, y=-0.4429782032966614, z=0.014744692512349077, roll=7.884650441866775e-28, pitch=-7.554679105908491e-28, yaw=2.245637386214381)
156table: Pose=RavenPose(x=0.0, y=-0.5, z=0.0, roll=0.0, pitch=-0.0, yaw=0.0)
@@
157Task goal: serve the banana inside the blue thing after drying it by placing on the plate
158
159#define assistant
160‘‘‘python
161def gen_plan(initial:RavenState):
162 plan = []
163 place_pose = PlaceSampler(initial, env, "plate").sample(rng)
164 banana_grasp = GraspSampler(initial, env, "banana").sample(rng)
165 plan.append(Action("pick", [[banana], grasp]))
@@
166 plan.append(Action("place_ontop", [[banana, plate], place_pose]))
167 bowl_bounds = get_aabb_bounds(initial, env, "bowl")
168 place_pose = PlaceSampler(initial, env, "bowl").sample(rng)
169 bowl_center_pose = get_obj_center(initial, env)
170 # Drop the object in the exact center of the bowl
171 place_pose[0] = bowl_center_pose[0]
172 place_pose[1] = bowl_center_pose[1]
173 place_pose[2] = bowl_center_pose + 0.05
@@
174 # Make sure to pick an orientation such that the banana will fit!
175 place_pose.roll = 1.5
176 banana_grasp = GraspSampler(initial, env, "banana").sample(rng)
177 plan.append(Action("pick", [[banana], grasp]))
@@
178 plan.append(Action("place_inside", [[banana, bowl], place_pose]))
179 return plan
180‘‘‘
181
182Here is yet another example for a different task:
183mug: Pose=RavenPose(x=-0.09269248694181442, y=-0.7042829990386963, z=0.026169249787926674, roll=0.0, pitch=-0.0, yaw=0.8605557025412023)
184potted_meat_can: Pose=RavenPose(x=0.17416073374449514, y=-0.33348321026557554, z=0.02017684663429707, roll=5.081222700168695e-05, pitch=0.00013538346655467005, yaw=-3.0371082921616765)
185plate: Pose=RavenPose(x=-0.11636300384998322, y=-0.4429782032966614, z=0.014744692512349077, roll=7.884650441866775e-28, pitch=-7.554679105908491e-28, yaw=2.245637386214381)
186table: Pose=RavenPose(x=0.0, y=-0.5, z=0.0, roll=0.0, pitch=-0.0, yaw=0.0)
187Task goal: serve spam from its can into the cup
188
189#define assistant
190‘‘‘python
191def gen_plan(initial:RavenState):
192 plan = []
193 spam_grasp = GraspSampler(initial, env, "potted_meat_can").sample(rng)
194 pour_params = PourSampler(initial, env, "potted_meat_can", "cup").sample(rng)
195 pour_params[3] = 2.0 # necessary for the contents to really fall into the cup.
@@
196 plan.append(Action("pick", [[potted_meat_can], spam_grasp]))
197 plan.append(Action("pour", [[potted_meat_can, mug], pour_params]))
198 return plan
199‘‘‘
200
201
@@
202Make sure to enclose your output with ‘‘‘python <output gen_plan(initial:Ravenstate): function> ‘‘‘ (ignore the angle brackets - those are just for illustrative purposes).’

### v4

@@ A.8 Code as Policies Full Prompt
You are operating in an environment that has access to the following classes: {env_structs}.
@@
You have access to the following variables. Note that TABLE_AABB.lower and TABLE_AABB.upper are tuples of length 3 that represent the lower and upper bounds on the x, y, and z positions that are on the table.
{typed_variables}
You also have access to the following objects in the environment.
Each of these has type RavenObject‘. You are not able to use other objects:
{object_names}
@@
Note that when you use/invoke these objects, do not use them as a string (i.e., do not provide quotes ’’). Instead
@@
use them directly (e.g. use banana, not banana’). Use the name of the object as a string only where explicitly required
(e.g. by the GraspSampler).
@@
Heres the poses of all the objects in the initial state of the scene (depicted in an attached image). Note that the axes of each object is shown as red (x-axis), green (y-axis), and blue (z-axis).
@@
Note that roll is rotation about the x-axis, pitch is rotation about the y-axis and yaw is rotation about the z-axis. Pay careful attention to the axes and the current orientations of objects in the initial state as and when you decide to write functions involving orientations.
{init_state}
@@
You also have access to helper functions whose signatures and docstrings are shown below. Pay careful attention to the arguments and return values of each function.
Helper function signatures:
{helper_functions_and_docstrings}
@@
You have access to the following set of skills expressed as pddl predicates followed by descriptions.
@@
You have no other skills you can use, and you must exactly follow the number of inputs described below.
@@
The coordinate axes are x, y, z where x is distance from the robot base, y is left/right from the robot base, and z is the height off the table.
Action("pick", [[o], g])
@@
Pick up object o at grasp g sampled from a grasp sampler. Note that you should use the object o directly and not use its name (e.g. banana and not banana’).
Action("place_ontop", [[o, s], p])
@@
If holding an object o (e.g. banana‘) place the object ontop surface s (e.g. table‘) at pose p. Note that you should invoke s and o as objects directly and not use their names (e.g. banana and not banana’).
Action("place_inside", [[o, c], p])
@@
If holding an object o (e.g. banana‘) place the object inside container c (e.g. bowl‘) at placement pose p. Note that you should invoke o and c as objects directly and not use their names (e.g. banana and not banana’).
Action("pour", [[o, c], p])
@@
If holding an object o (e.g. tomato_soup_can‘) pour from the object onto/inside of container c (e.g. bowl‘) at a pouring pose dictated by p. Specifically, p is 4 numbers: <x, y, z> corresponding to the position to reach, and <theta> corresponding to the angle (in radians; -3.14 to 3.14) to tilt the hand to accomplish the pour. Note that you should invoke o and c as objects directly and not use their names (e.g. banana and not banana’).
@@
Finally, to help you generate continuous parameters required for these skills, you have access to the following samplers.
@dataclass
class GraspSampler(Sampler):
curr_state: RavenState
env: RavenYCBEnv
object_name: str
def sample(self, rng: np.random.Generator) -> List[float]:
"""Simply return some value within the AABB of the object, and at any
orientation."""
aabb_lb, aabb_ub = get_aabb_bounds(
self.curr_state, self.env, self.object_name
lb = (aabb_lb[0], aabb_lb[1], aabb_lb[2], -3.14159265, -3.14159265, -3.14159265)
ub = (aabb_ub[0], aabb_ub[1], aabb_ub[2], 3.14159265, 3.14159265, 3.14159265)
return rng.uniform(lb, ub).tolist()
@dataclass
class PlaceSampler(Sampler):
curr_state: RavenState
env: RavenYCBEnv
object_name_to_place_on_or_inside: str
def sample(self, rng: np.random.Generator) -> List[float]:
"""Simply return some value within the AABB of the object, and at any
orientation."""
min_drop_height = 0.01
max_drop_height = 0.35
aabb_lb, aabb_ub = get_aabb_bounds(
self.curr_state, self.env,
self.object_name_to_ place_on_or_inside
lb = (
aabb_lb[0],
aabb_lb[1],
aabb_lb[2] + min_drop_height,
-3.14159265,
-3.14159265,
-3.14159265,
ub = (
aabb_ub[0],
aabb_ub[1],
aabb_ub[2] + max_drop_height,
3.14159265,
3.14159265,
3.14159265,
return rng.uniform(lb, ub).tolist()
@dataclass
class PourSampler(Sampler):
curr_state: RavenState
env: RavenYCBEnv
holding_obj_name: str
obj_name_to_pour_into_or_ontop: str
def sample(self, rng: np.random.Generator) -> List[float]:
@@
"""Return a position sampled from above obj_name_to_pour_ into_or_ontop
and a pitch randomly sampled between bounds."""
# We sample (x, y, z, pitch.)
obj_holding_lb, obj_holding_ub = get_aabb_bounds(
self.curr_state, self.env, self.holding_obj_name
min_pour_height = obj_holding_ub[2] - obj_holding_lb[2]
max_pour_height = min_pour_height * 2
aabb_lb, aabb_ub = get_aabb_bounds(
self.curr_state, self.env, self.obj_name_to_pour _into_or_ontop
pitch_range = (-3.14, 3.14)
lb = (aabb_lb[0], aabb_lb[1], aabb_lb[2] + min_pour_height, pitch_range[0])
ub = (aabb_ub[0], aabb_ub[1], aabb_ub[2] + max_pour_height, pitch_range[1])
return rng.uniform(lb, ub).tolist()
@@
Your goal is to generate a python function that returns a plan that performs the provided task. This function can
@@
use helper functions that must be defined within the scope of the function itself.
@@
The main function should be named EXACTLY gen_plan, and it should take in only one parameter corresponding to the environment state as input. Do not change the names. Do not create any additional classes or overwrite any existing ones. You are only allowed to create helper functions inside the gen_plan function.
Current provided goal: {provided_task_goal}
Here is an example gen_plan function for a different task:
#define user
Init state:
bowl: Pose=RavenPose(x=-0.09269248694181442, y=-0.7042829990386963, z=0.026169249787926674, roll=0.0, pitch=-0.0, yaw=0.8605557025412023)
banana: Pose=RavenPose(x=0.17416073374449514, y=-0.33348321026557554, z=0.02017684663429707, roll=5.081222700168695e-05, pitch=0.00013538346655467005, yaw=-3.0371082921616765)
plate: Pose=RavenPose(x=-0.11636300384998322, y=-0.4429782032966614, z=0.014744692512349077, roll=7.884650441866775e-28, pitch=-7.554679105908491e-28, yaw=2.245637386214381)
table: Pose=RavenPose(x=0.0, y=-0.5, z=0.0, roll=0.0, pitch=-0.0, yaw=0.0)
Task goal: put the lemon on the plate and the banana on the table
#define assistant
‘‘‘python
def gen_plan(initial:RavenState):
plan = []
place_pose = PlaceSampler(initial, env, "plate").sample(rng)
lemon_grasp = GraspSampler(initial, env, "lemon").sample(rng)
plan.append(Action("pick", [[lemon], lemon_grasp]))
@@
plan.append(Action("place_ontop", [[lemon, plate], place_pose]))
place_pose = PlaceSampler(initial, env, "table").sample(rng)
banana_grasp = GraspSampler(initial, env, "banana").sample(rng)
plan.append(Action("pick", [[banana], banana_grasp]))
@@
plan.append(Action("place_ontop", [[banana, table], RavenPose(x=x_place, y=y_place, z=0.02)]))
return plan
@@
Here is another example for a different goal involving the same objects and initial state. Note that in this case, the bowl is blue.
@@
Also note that the banana needs to be rotated so that it fits into the bowl; in general you should pay careful
@@
attention to any angular constraints that might be important for solving different tasks.
#define user
Init state:
bowl: Pose=RavenPose(x=-0.09269248694181442, y=-0.7042829990386963, z=0.026169249787926674, roll=0.0, pitch=-0.0, yaw=0.8605557025412023)
banana: Pose=RavenPose(x=0.17416073374449514, y=-0.33348321026557554, z=0.02017684663429707, roll=5.081222700168695e-05, pitch=0.00013538346655467005, yaw=-3.0371082921616765)
plate: Pose=RavenPose(x=-0.11636300384998322, y=-0.4429782032966614, z=0.014744692512349077, roll=7.884650441866775e-28, pitch=-7.554679105908491e-28, yaw=2.245637386214381)
table: Pose=RavenPose(x=0.0, y=-0.5, z=0.0, roll=0.0, pitch=-0.0, yaw=0.0)
@@
Task goal: serve the banana inside the blue thing after drying it by placing on the plate
#define assistant
‘‘‘python
def gen_plan(initial:RavenState):
plan = []
place_pose = PlaceSampler(initial, env, "plate").sample(rng)
banana_grasp = GraspSampler(initial, env, "banana").sample(rng)
plan.append(Action("pick", [[banana], grasp]))
@@
plan.append(Action("place_ontop", [[banana, plate], place_pose]))
bowl_bounds = get_aabb_bounds(initial, env, "bowl")
place_pose = PlaceSampler(initial, env, "bowl").sample(rng)
bowl_center_pose = get_obj_center(initial, env)
# Drop the object in the exact center of the bowl
place_pose[0] = bowl_center_pose[0]
place_pose[1] = bowl_center_pose[1]
place_pose[2] = bowl_center_pose + 0.05
@@
# Make sure to pick an orientation such that the banana will fit!
place_pose.roll = 1.5
banana_grasp = GraspSampler(initial, env, "banana").sample(rng)
plan.append(Action("pick", [[banana], grasp]))
@@
plan.append(Action("place_inside", [[banana, bowl], place_pose]))
return plan
Here is yet another example for a different task:
mug: Pose=RavenPose(x=-0.09269248694181442, y=-0.7042829990386963, z=0.026169249787926674, roll=0.0, pitch=-0.0, yaw=0.8605557025412023)
potted_meat_can: Pose=RavenPose(x=0.17416073374449514, y=-0.33348321026557554, z=0.02017684663429707, roll=5.081222700168695e-05, pitch=0.00013538346655467005, yaw=-3.0371082921616765)
plate: Pose=RavenPose(x=-0.11636300384998322, y=-0.4429782032966614, z=0.014744692512349077, roll=7.884650441866775e-28, pitch=-7.554679105908491e-28, yaw=2.245637386214381)
table: Pose=RavenPose(x=0.0, y=-0.5, z=0.0, roll=0.0, pitch=-0.0, yaw=0.0)
Task goal: serve spam from its can into the cup
#define assistant
‘‘‘python
def gen_plan(initial:RavenState):
plan = []
spam_grasp = GraspSampler(initial, env, "potted_meat_can").sample(rng)
pour_params = PourSampler(initial, env, "potted_meat_can", "cup").sample(rng)
pour_params[3] = 2.0 # necessary for the contents to really fall into the cup.
@@
plan.append(Action("pick", [[potted_meat_can], spam_grasp]))
plan.append(Action("pour", [[potted_meat_can, mug], pour_params]))
return plan
@@
Make sure to enclose your output with ‘‘‘python <output gen_plan(initial:Ravenstate): function> ‘‘‘ (ignore the angle brackets - those are just for illustrative purposes).